OpenStudy (anonymous):

proportion: 6x-5/6=x/8 (the '/' means a fraction bar)

OpenStudy (anonymous):

\[\frac{6x-5}{6}=\frac{x}{8}\]some people remember how to solve this type as a "cross multiplication"\[8(6x-5)=6x\]distribute\[48x-40=6x\]transpose the 40 and the 6x\[48x-6x=40\]combine like terms\[42x=40\]divide by 42 and reduce\[x=\frac{40}{42}=\frac{10}{21}\]

OpenStudy (anonymous):

sorry! numerator is supposed to be 20 NOT 10!!!!!!!!!!!!!!!!!!!!!!!!!!!!